# What could go wrong? Discovering and describing failure modes in computer vision

### v1

@@ 6 Experiments
Table 1: Examples of sentences from the user-defined sentence set ${\mathcal{S}}$. For each sentence we show if it belongs to ${\mathcal{S}}^{*}_{\beta}$ and to ${\mathcal{R}}_{\mathcal{S}}$ for TopS (TS), PDiff (PD), SetDiff (SD) and FPDiff (FP) for the datasets WD2, IDD and ACDC. “” means $s_{n}\in{\mathcal{S}}^{*}_{\beta}$ and “$\times$” means $s_{n}\notin{\mathcal{S}}^{*}_{\beta}$—namely, the hardness score of $s_{n}$ is below the required level. For the methods, “+”/“+” indicate that the sentence is in ${\mathcal{R}}_{\mathcal{S}}$, where “+” means true positive (i.e., the sentence is also in ${\mathcal{S}}^{*}_{\beta}$) while “+” means false positive. An empty space indicates that the sentence was not in the corresponding ${\mathcal{R}}_{\mathcal{S}}$.
@@
Default design choices. We use Open-CLIP [17] as vision-and-language embedding space. To generate the GT sentence set ${\mathcal{S}}^{*}_{\beta}$, we use $\beta=.2*\omega_{\theta}^{\text{std}}$ with $\omega_{\theta}^{\text{std}}$ being the standard deviation of $\omega_{\theta}$ computed over ${\mathcal{X}}$. We set the number of hard clusters to $C=15$ for large datasets and $C=5$ for the per-class data analyses. We use the same number of clusters for the easy set ($|calC^{h}|=|{\mathcal{C}}^{e}|=C$). To split the data we use the output entropy in the unsupervised cases and class probability or ranking in the per-class case (see details in Appendix D). We retain three sentences for each cluster($K=3$).
@@
Figure 3: Two hard clusters in ACDC explained by different methods. We illustrates a few example images from the cluster and below the three sentences retained by different methods for this cluster.
@@ 6.1 Results
In Figs. 3, 5, 7, 9 and 11 we report qualitative results, showing the sentences selected by our methods for various hard clusters. In all cases, each row represents a hard cluster, and below we report the three sentences provided by the different methods.
Focusing, e.g., on Figs. 3, 5 and 7 (urban segmentation on ACDC, IDD and WD2, respectively) we can see that our methods can successfully isolate the main reasons for clusters to be hard, for example difficult visual conditions (rainy/foggy weather or image taken at night) or the presence of elements unseen during training such as tunnel,mud or light reflections). Indeed, such elements can represent real challenges for a model trained on Cityscapes as observed also e.g. in [4].
With our methods, the user can easily asses the potential failures at a glance, instead of inspecting the hundreds of images contained in the above clusters [^4] The user could rank easily the images in the cluster for each selected sentence for the visual inspection based on the corresponding Open-CLIP similarity scores.. Such information can be further used for different down-stream tasks, e.g., helping the user to select which new samples to annotate [41, 33].
In Fig. 2, we provide numerical results for the different methods, tasks and datasets, using the metrics introduced in Sec. 5 and the default setting and hyper-parameters detailed above. First, we observe that TopS  is the best performing method in terms of ACR scores. This is not surprising, since sentences closest to the prototype have obviously the highest coverage. Yet, these sentences are not necessarily pointing to error explanations—indeed, see the low AHR scores in some cases, with IDD being the most dramatic one (right-most column, third panel).
Concerning the metrics related to the failure explanations (AHR/TPR/JI), in both NICO++{++}+ +  scenarios, where we have ground truth relevance scores, and on WD2 FPDiff performs best or close to it. It performs worse on ACDC and IDD: there is no clear winner among the proposed methods in these two datasets, since the higher performance of PDiff  and SetDiff  in terms of TPR and JI comes at the price of a lower ACR.
@@
Figure 5: Two hard clusters from IDD explained by different methods. We illustrates a few example images from the cluster and below the three sentences retained by different methods for this cluster.
@@ 6.1.1 ACDC
To recap, we test models trained on Cityscapes (daylight and clear weather) on ACDC, which contains samples recorded under various weather and daylight conditions. Results in Fig. 2 show that the best AHR performance is obtained with TopS , but also that this method is the worst performing one in terms of TPR and JI. For what concerns the latter metrics, PDiff  shines as the best performing method. To shed light on these results, we complement them with Tab. 1 (and with the more detailed Tab. 5) showing sentences from the User-defined sentence set ${\mathcal{S}}$ that are either in ${\mathcal{S}}^{*}_{\beta}$ or in any ${\mathcal{R}}_{\mathcal{S}}$, namely the ground truth set of hard sentences and the ones retrieved by our methods, respectively.
@@
We can observe that while using the GT mIoU improved significantly the ACR and AHR scores (averaged over all clusters), globally the TPR and JI changed only slightly. This suggests that, in general, the sentences retained by PDiff , FPDiff and SetDiff  have better coverage in the cluster (higher ACR) and retain more often a hardness score that satisfies the required constraint. The TPR and JI scores are comparable across splitting strategy. By analyzing the sentences retrieved in each setting, we discovered that with GT mIoU the contrastive methods are able to recover sentences such as “taken at night”, “taken in the evening” and “taken at dusk”, but fail in retrieving sentences such as “with road barrier”/“with rail track on the road” and “with motion blur“/”with underexposure”, sentences that were retrieved by using the entropy.
Finally, sentences referring to fog and rain, having scores slightly below the required condition $\omega_{\theta}^{s_{n}}<\omega_{\theta}^{\text{avg}}-\beta$, are present in ${\mathcal{R}}_{\mathcal{S}}$ for all splitting strategies and most methods. One could interpret those as less severe false positives.
@@ 6.1.2 IDD
6.1.3 WD2
GT-based UserDef LLM-based
@@ 6.1.3 WD2
GT-based UserDef LLM-based
@@
GT-based UserDef LLM-based
@@
Varying the number of retained sentences ($K$). We show in Fig. 6 (top) results for IDD as we vary the number of retained sentences per cluster. We observe that with the increase of the number of sentences retained, the ACR increases for FPDiff , but decreases for the other methods; on the other hand, the ACR is stable starting from $K=5$. The AHR is generally stable as we vary $K$, except for FPDiff , where the best value is obtained with $K=1$. Both TPR and JI benefit from the increase of the number of sentences (higher $K$). Overall, we can confirm that SetDiff  outperforms the other methods in terms of AHR, TPR and JI. This result, combined with the low ACR, suggests that SetDiff  is more capable of detecting rarer failure reasons, for which there are less support example images in the cluster/dataset.
Varying the number of clusters ($C$). In Fig. 6 (bottom) we present results obtained on IDD when we vary the number of clusters ($|{\mathcal{C}}^{h}|=|{\mathcal{C}}^{e}|=C$). By increasing the number of clusters, we observe an increase of the TPR (as expected) for all methods—with PDiff  being the one that benefits the most. Concerning the other metrics, the default value $C=15$ is a reasonable trade off. These experiments suggest that IDD is overall a very challenging set, with some of the failure causes under-represented—especially the ones related to the weather conditions (see sentences in Tab. 1 and Tab. 5), or not clustered together according to these characteristics—cf. Fig. 5. Furthermore, note that as the filtering in FPDiff relies on the TopS  ranking, hence, the very poor performance of TopS  on this particular set (IDD) affects negatively the performance of FPDiff , making PDiff  perform better than FPDiff when we consider more clusters or retain more sentences. Finally, the poor performance obtained with $C=1$—corresponding to running the methods without the clustering step—show that describing the full hard set by contrasting with the full easy set is a sub-optimal solution.
@@
Figure 9: Three out of 15 hard clusters from NICO++{++}+ + 85 (unsupervised case) explained by different methods.
Sensitivity to the choice of the sentence set. We analyse here the performance of the different methods as we vary the sentence set ${\mathcal{S}}$. For this experiment, we use the WD2 dataset, because for this dataset we have both User-defined and GT-based sentence sets. We also compare with the significantly larger LLM-based sentence set (see details in Appendix C). We show qualitative comparison in Fig. 7 and numerical results in Fig. 8.
First, we notice that both the User-defined and the LLM-generated sentence sets yield similar ACR and AHR scores. Yet, for the latter we have many sentences that are semantically related and carry similar hardness scores, for example “taken at night”/“mesmerizing moonlight”/“spellbound starlight”/“enigmatic night”/“miraculous midnight”. Since we limit the number of sentences each method can retrieve, in this case we cannot retrieve all of them with a small $K$ and, in turn, this will result into significantly lower TPR and JI scores.
When we use the GT-based sentence set with $o=0.2$, ${\mathcal{S}}^{*}_{\beta}$ contains only five sentences that are “showing a tunnel”, “showing a bridge”, “showing traffic light”, “showing a bus” and “showing snow”. Here, all methods introduce many false positives (explaining the low JI score). PDiff  and SetDiff  were able to retain four out of five sentences (except “showing a bus”) and TopS  and FPDiff  three out of five (missing also “showing a bridge”). The missing sentences had not enough image coverage in the clusters. The AHR and ACR scores are also low, due to the fact that it is not easy to find shared content for all clusters (see the low ACR even for TopS ). Finally, note that the GT-based sentence set mainly describes the presence of the classes (“Image showing <class>"), therefore, it is not sufficient to describe the hard clusters to the desired level of details. This highlights the importance of properly designing ${\mathcal{S}}$, as well as the importance of using additional text resources to analyze the performance of our models.
Figure 10: Varying $\beta$ (by varying $o$). Results obtained on NICO++{++}+ + 85 (unsupervised). The value $\beta$ defines the ground truth set ${\mathcal{S}}^{*}$, so its value impacts only the AHR/TPR/JI metrics (not the output of the methods or ACR). On x-axis: we plot $o$, where $\beta=o\cdot\omega_{\theta}^{\text{std}}$ with $\omega_{\theta}^{\text{std}}$ being the standard deviation of $\omega_{\theta}$ (entropy) computed over ${\mathcal{X}}$.
@@
Figure 11: Analyzing EfficientNet performance on the ImageNet 1K. The first two rows are from the “acoustic guitar” class, and the third is from the bath tube class.
6.1.4 NICO++{++}+ + 85
We provide in LABEL:{fig:qualitativeNPPmain} qualitative examples for NICO++{++}+ + 85, in the unsupervised case (where we use the prediction entropy to split the data). We observe that our methods can capture the “context-class” associations that were rarely seen in the training set [^5] See in Tab. 2 the distribution of the images in the validation set over the different classes-context pairs. Note that the validation set follows the training distribution. and hence the model struggles with their classification. Indeed we find “birds” or “plants” in “water” (top row), “landways” in the “rocks” (middle row) and “landways” in ‘dim lighting” (bottom row).
@@ 6.1.4 NICO++{++}+ + 85
Varying $\beta$ (by varying $o$). Recall that $\beta$ is the hyper-parameter used to select the ground-truth sentence set ${\mathcal{S}}^{*}$, where higher $\beta$ values imply retaining only the sentences with the highest hardness scores as valid failure descriptions, while lower values imply that sentences with a score $\omega_{\theta}^{s_{n}}$ below $\omega_{\theta}^{\text{avg}}$ but with a smaller margin $\beta$ are also considered as valid explanations. In Fig. 10 we show the impact of varying $\beta$ on the AHR, TPR, and JI metrics. We carried out this experiment on the spurious classification with NICO++{++}+ + 85, considering the unsupervised case where we have GT image-sentence relevance scores (top row). As the number of sentences retained by the method is fixed ($K=3$), varying $\beta$ only affects ${\mathcal{S}}^{*}$ and not ${\mathcal{R}}_{\mathcal{S}}$. Therefore, the value of ACR (that does not depend on $\beta$) is the same for each $o$ value. Intuitively, with this study we assess how much each method focuses on the hardest sentences. Indeed, increasing the value of $o$ (and hence increasing $\beta$) corresponds to accepting fewer and fewer sentences as valid explanations, namely the ones with the highest hardness score (lowest average accuracy). We observe that the FPDiff method, while also experiencing a drop in performance as we make the problem harder, holds a better performance than the others methods for all the values of $o$.
@@ 6.1.5 ImageNet 1K
Finally, we focus on ImageNet 1K, where we treat each class separately. We can make several observations. i) When the error slices (hard clusters) mainly contain false positives from semantically similar classes or false negatives with shared content, TopS will generally capture the shared characteristics—see examples in Fig. 11 (top two rows), Fig. 13 and Fig. 14 in Appendix E. ii) In the case of false positives with semantically similar classes, e.g.electric guitar” and “banjo” in place of the “acoustic guitar” class (see first row in Fig. 11), the main difficulty for the other methods is that the same sentences are also highly ranked for the easy clusters, making it difficult for them to correctly identify the desired explanations—namely, contrasting the explanation between the easy and the hard cluster. This limitation is inherited from CLIP, which has difficulties to distinguish fine-grain classes. iii) When the false negatives share some semantic similarities and there is a candidate sentence that well describes this, all methods are able to find it. For example, see in Fig. 11 (second row) the difficulty of the model to label the “acoustic guitar” when the photo is showing the person holding the guitar on a “stage”. Note that for these images PDiff and FPDiff also identified the blur as potential challenge (or “low contrast” and “dim weather” condition, in Fig. 13). iv) Finally, we have clusters where even for a human is difficult to describe what is common in the images beyond the object class. When there is a mix between false positives and false negatives, all methods including TopS have difficulties to identify the failure reason. In these cases, the selected sentences are relevant to some of the images in the cluster but they do not point to the reason for the failure (see Fig. 11 last row).
@@ 7 Concluding remarks
In the future, we plan to overcome the need to provide the sentence set ${\mathcal{S}}$, a design choice that makes evaluation more tractable, but that also limits our methods. Defining the sentence set ${\mathcal{S}}$ a priori has the advantage that our methods can be deployed off the shelf without any overhead, but yet the set might omit certain unexpected failure reasons. A possible solution could be to build the sentence set ${\mathcal{S}}$ directly from the target set ${\mathcal{X}}$, e.g., by running a captioner to describe all its images and merge these descriptions into the sentence set. Yet, in addition to being prohibitively expensive, a solution like this needs to be iterated for every new target set before running any method. Furthermore, in the case of a large image set, it would also require further steps to remove redundancy and to regroup similar sentences.

### v2

@@ 6 Experiments
Table 1: Examples of sentences from the User-defined sentence set ${\mathcal{S}}$. For each sentence we show if it belongs to ${\mathcal{S}}^{*}_{\beta}$ and to ${\mathcal{R}}_{\mathcal{S}}$ for TopS (TS), PDiff (PD), SetDiff (SD) and FPDiff (FP) for the datasets WD2, IDD and ACDC. “” means $s_{n}\in{\mathcal{S}}^{*}_{\beta}$ and “$\times$” means $s_{n}\notin{\mathcal{S}}^{*}_{\beta}$—namely, the hardness score of $s_{n}$ is below the required level. For the methods, “+”/“+” indicate that the sentence is in ${\mathcal{R}}_{\mathcal{S}}$, where “+” means true positive (i.e., the sentence is also in ${\mathcal{S}}^{*}_{\beta}$) while “+” means false positive. An empty space indicates that the sentence is not in the corresponding ${\mathcal{R}}_{\mathcal{S}}$.
@@
Default design choices. We use Open-CLIP [17] as vision-and-language embedding space. ${\mathcal{S}}$ is a User-defined sentence set in the case of urban scene segmentation and ImageNet and GT-based for NICO++{++}+ + (see details in Sec. C.1). To generate the ground-truth sentence set ${\mathcal{S}}^{*}_{\beta}$ containing the sentences with valid hardness score (see Sec. C.2), we set $\beta=.2*\omega_{\theta}^{\text{std}}$ with $\omega_{\theta}^{\text{std}}$ being the standard deviation of $\omega_{\theta}$ computed over ${\mathcal{X}}$. We set the number of hard clusters to $C=15$ for large datasets and $C=5$ for the per-class data analyses. We use the same number of clusters for the easy set ($|{\mathcal{C}}^{h}|=|{\mathcal{C}}^{e}|=C$). To split the data we use the output entropy in the unsupervised cases and class probability or ranking in the per-class case (see details in Appendix D). We retain three sentences for each cluster ($K=3$).
@@
Figure 3: Two hard clusters in ACDC explained by different methods. We illustrates a few example images from the cluster and below the three sentences retained by different methods for it.
@@ 6.1 Results
In Figs. 3, 5, 7, 9 and 11 we report qualitative results, showing the sentences selected by our methods for various hard clusters. In all cases, each row represents a hard cluster, and at the bottom we report the three sentences provided by the different methods.
Focusing, e.g., on Figs. 3, 5 and 7 (urban segmentation on ACDC, IDD and WD2, respectively) we can see that our methods can successfully isolate the main reasons for the clusters to be hard, for example difficult visual conditions (rainy/foggy weather or image taken at night) or the presence of elements unseen during training such as tunnels, mud or light reflections). Indeed, such elements can represent real challenges for a model trained on Cityscapes, as observed also, e.g., in [4].
With our methods, the user can easily asses the potential failures at a glance, instead of inspecting the hundreds of images contained in the above clusters. They could easily rank the images in the cluster for each selected sentence for the visual inspection based on the corresponding Open-CLIP similarity scores. Such information can be further used for different down-stream tasks, for example helping the user to select which new samples to annotate [41, 33].
In Fig. 2, we provide numerical results for the different methods, tasks and datasets, using the metrics introduced in Sec. 5 and the default setting and hyper-parameters detailed above. First, we observe that TopS  is the best performing method in terms of ACR scores. This is not surprising, since sentences closest to the prototype have obviously the highest coverage. Yet, these sentences are not necessarily pointing to error explanations—indeed, see the low AHR scores in some cases, with IDD being the most dramatic one (right-most column, third panel). Concerning the metrics related to the failure explanations (AHR/TPR/JI), in both NICO++{++}+ +  scenarios, where we have ground truth relevance scores, and on WD2 FPDiff performs best or close to it. It performs worse on ACDC and IDD: there is no clear winner among the proposed methods in these two datasets, since the higher performance of PDiff  and SetDiff  in terms of TPR and JI comes at the price of a lower ACR.
@@
Figure 5: Two hard clusters from IDD explained by different methods. We illustrates a few example images from the cluster and below the three sentences retained by different methods for it.
@@ 6.1.1 ACDC
To recap, we test models trained on Cityscapes (daylight and clear weather) on ACDC, which contains samples recorded under various weather and daylight conditions. Results in Fig. 2 show that the best AHR performance is obtained with TopS , but also that this method is the worst performing one in terms of TPR and JI. For what concerns the latter metrics, PDiff  shines as the best performing method. To shed light on these results, we complement them with Tab. 1 (and with the more detailed Tab. 5) showing sentences from the sentence set ${\mathcal{S}}$ that are either in ${\mathcal{S}}^{*}_{\beta}$ or in any ${\mathcal{R}}_{\mathcal{S}}$, namely the ground truth set of hard sentences and the ones retrieved by our methods, respectively.
@@
We can observe that while using the GT mIoU improved significantly the ACR and AHR scores (averaged over all clusters), globally the TPR and JI changed only slightly. This suggests that, in general, the sentences retained by PDiff , FPDiff and SetDiff  have better coverage in the cluster (higher ACR) and retain more often a hardness score that satisfies the required constraint. The TPR and JI scores are comparable across splitting strategy. By analyzing the sentences retrieved in each setting, we discovered that with GT mIoU the contrastive methods are able to recover sentences such as “taken at night”, “taken in the evening” and “taken at dusk”, but fail in retrieving sentences such as “with road barrier”/“with rail track on the road” and “with motion blur“/”with underexposure”, namely sentences that were retrieved by using the entropy.
Finally, sentences referring to fog and rain, having scores slightly below the required condition ($\omega_{\theta}^{s_{n}}<\omega_{\theta}^{\text{avg}}-\beta$), are present in ${\mathcal{R}}_{\mathcal{S}}$ for all splitting strategies and most methods. One could interpret those as less severe false positives.
@@ 6.1.2 IDD
GT-based User-defined LLM-based
@@
GT-based User-defined LLM-based
@@
GT-based User-defined LLM-based
@@
Varying the number of retained sentences ($K$). We show in Fig. 6 (top) results for IDD as we vary the number of retained sentences per cluster. By increasing the number of sentences retained, we observe that the ACR increases for FPDiff , but decreases for the other methods; on the other hand, the ACR is stable starting from $K=5$. The AHR is generally stable as we vary $K$, except for FPDiff , where we obtain the best value with $K=1$. Both TPR and JI benefit from the increase of the number of sentences (higher $K$). Overall, SetDiff  outperforms the other methods in terms of AHR, TPR and JI. This result, combined with the low ACR, suggests that SetDiff  is more capable of detecting rarer failure reasons, for which there are less support example images in the cluster/dataset.
Varying the number of clusters ($C$). In Fig. 6 (bottom) we present results obtained on IDD when we vary the number of clusters ($|{\mathcal{C}}^{h}|=|{\mathcal{C}}^{e}|=C$). By increasing the number of clusters, we observe an increase of the TPR (as expected) for all methods—with PDiff  being the one that benefits the most. Concerning the other metrics, the default value $C=15$ is a reasonable trade off. These experiments suggest that IDD is overall a very challenging set, with some of the failure causes under-represented—especially the ones related to the weather conditions (see sentences in Tab. 1 and Tab. 5), or not clustered together according to these characteristics (cf. Fig. 5). Furthermore, note that as the filtering in FPDiff relies on the TopS  ranking, the very poor performance of the latter on this particular set (IDD) affects negatively the performance of the former, making PDiff  perform better than FPDiff when we consider more clusters or retain more sentences. Finally, the low performance obtained with $C=1$—corresponding to running the methods without the clustering step—shows that describing the full hard set by contrasting with the full easy set is a sub-optimal solution.
@@
[…]
6.1.3 WD2
We analyse the performance of the different methods on WD2, varying the sentence set ${\mathcal{S}}$. Indeed, for this dataset we have both User-defined (the same used throughout the experimental validation), a GT-based sentence sets, as well as a significantly larger LLM-based sentence set (see Appendix C). We show qualitative comparison in Fig. 7 and numerical results in Fig. 8.
Sensitivity to the choice of the sentence set. First, we notice that both the User-defined and the LLM-generated sentence sets yield similar ACR and AHR scores. Yet, for the latter we have many sentences that are semantically related and carry similar hardness scores, for example “taken at night”/“mesmerizing moonlight”/“spellbound starlight”/“enigmatic night”/“miraculous midnight”. Since we limit the number of sentences each method can retrieve, in this case we cannot retrieve all of them with a small $K$ and, in turn, this will result into significantly lower TPR and JI scores.
When we use the GT-based sentence set with $o=0.2$, ${\mathcal{S}}^{*}_{\beta}$ contains only five sentences that are “showing a tunnel”, “showing a bridge”, “showing traffic light”, “showing a bus” and “showing snow”. Here, all methods introduce many false positives (explaining the low JI score). PDiff  and SetDiff  are able to retain four out of five sentences (except “showing a bus”) and TopS  and FPDiff  three out of five (missing also “showing a bridge”). The missing sentences had not enough image coverage in the clusters. The AHR and ACR scores are also low, due to the fact that it is not easy to find shared content for all clusters (see the low ACR even for TopS ). Finally, note that the GT-based sentence set mainly describes the presence of the classes (“Image showing <class>”), therefore, it is not sufficient to describe the hard clusters to the desired level of details. This highlights the importance of properly designing ${\mathcal{S}}$, as well as the importance of using additional textual metadata to analyze the performance of our models.
[…]
@@ 6.1.3 WD2
Figure 11: Analyzing EfficientNet performance on the ImageNet 1K. The first two rows are from the “acoustic guitar” class, and the third is from the “bath tube” class.
[…]
@@
[…]
@@ 6.1.5 ImageNet 1K
Finally, we focus on ImageNet 1K, where we treat each class separately. We can make several observations. i) When the error slices (hard clusters) mainly contain false positives from semantically similar classes or false negatives with shared content, TopS will generally capture the shared characteristics—see examples in Fig. 11 (top two rows), Fig. 12 and Fig. 13 in Appendix E. ii) In the case of false positives with semantically similar classes, e.g., “electric guitar” and “banjo” in place of the “acoustic guitar” class (see first row in Fig. 11), the main difficulty for the other methods is that the same sentences are also highly ranked for the easy clusters, making it difficult for them to correctly identify the desired explanations—namely, contrasting the explanation between the easy and the hard clusters. This limitation is inherited from CLIP, which has difficulties to distinguish fine-grain classes. iii) When the false negatives share some semantic similarities and there is a candidate sentence that well describes this, all methods are able to find it. For example, see in Fig. 11 (second row) the difficulty of the model to label the “acoustic guitar” when the photo is showing the person holding the guitar on a “stage”. Note that for these images PDiff and FPDiff also identified the “blur” as potential challenge (or “low contrast” and “dim weather” condition, in Fig. 12). iv) Finally, we have clusters where even for a human is difficult to describe what is common in the images beyond the object class. When there is a mix between false positives and false negatives, all methods including TopS have difficulties to identify the failure reason. In these cases, the selected sentences are relevant to some of the images in the cluster but they do not point to the reason for the failure (see Fig. 11 last row).
@@ 7 Concluding remarks
In the future, we plan to overcome the need to define the sentence set ${\mathcal{S}}$ a priori. Indeed, while this design choice makes the evaluation more tractable and allows deploying our methods off the shelf, the set might omit certain unexpected failure reasons. A possible solution could be to build the sentence set ${\mathcal{S}}$ directly from the target set ${\mathcal{X}}$, e.g., by running a captioner to describe all its images and merge these descriptions into the sentence set. Yet, in addition to being prohibitively expensive, a solution like this needs to be iterated for every new target set before running any method. Furthermore, in the case of a large image set, it would also require further steps to remove redundancy and to regroup similar sentences.